you and your friend are walking around a tennis court that has a 10-foot high fence around it. You see two tennis balls lying outside the fence. You pick up a ball and try to throw it with an initial velocity of 22 feet per second. Did the ball made it over the fence?
Answer by Alan3354(69443)

you and your friend are walking around a tennis court that has a 10-foot high fence around it. You see two tennis balls lying outside the fence. You pick up a ball and try to throw it with an initial velocity of 22 feet per second. Did the ball made it over the fence?
------------
h(t) = -16t^2 + 22t, height in feet, t in seconds
----
The max ht is the vertex
t at the vertex = -b/2a = -22/-32 = 11/16 seconds
---
h(11/16) = -16(121/256) + 22*(11/16) (max ht)
= 242/16 - 121/16 ft
= 7.5625 feet
Not high enough